OpenStudy (jennychan12):

integral (1/(1+rad(2x))dx use u-sub what did they do here?

OpenStudy (jennychan12):

where did the u-1 come from?

OpenStudy (abb0t):

\[\int\limits \frac{ 1 }{ 1+\sqrt{2x} }dx???\]

OpenStudy (jennychan12):

yeah

OpenStudy (abb0t):

I would of used conjugates for this, but if you are using u-sub then \[u = 1+\sqrt{2x}\] \[(u-1)=\sqrt{2x}\] what they did there was just find a substitute. Since you don't have anything you can substitute for sqrt, they just rearranged it. So you would get:\[\int\limits \frac{ 1 }{ 1+(u-1) }du\]

OpenStudy (anonymous):

If you said u=sqrt(x), then du cant be = dx, i feel like the question is somewhat wrong lol

OpenStudy (abb0t):

The original problem might be off, but that's the same idea. I would of multiplied by conjugates.

OpenStudy (jennychan12):

*would have i would too, but that's what the directions say if u = 1-rad(2x) du = [(1/2)(2x)^-1/2(2])dx du = x^-1/2 rad(2x)du = dx oh wait. i think i get it now. -_-